Assume that s parameter x is described by a continuous function f. A mean value of parameter x is 24.2. The standard deviation of that parameter is 4.6%. Assume that parameter x is normally distributed. What is the probability of x to be between 26.5 and 26.73?
Answer #1

Solution :

Given that ,

mean = = 24.2

standard deviation = = 4.6

P( 26.5 < x < 26.73) = P[(26.5 - 24.2)/ 4.6 ) < (x - ) /  < (26.73 - 24.2) / 4.6) ]

= P(0.50 < z < 0.55)

= P(z < 0.55) - P(z < 0.50)

Using z table,

= 0.7088 - 0.6915

= 0.0173